OpenStudy (anonymous):

Explain why a matrix that does not have the same number of rows and columns cannot have a multiplicative inverse.

jimthompson5910 (jim_thompson5910):

If an m x n matrix A (where m doesn't equal n) had a multiplicative inverse, then A*A^(-1) = A^(-1)*A = I where I is the identity matrix But multiplication isn't defined because the number of columns in A is not the same as the number of rows in A^(-1)

jimthompson5910 (jim_thompson5910):

So A has to be square